A pharmaceutical company produces a very rare drug for cancer treatment but they have very limited resources for it & they will able to produce very less quantities so company wants to track each drug & its operations like (locations, user, shipping etc.) but with their current system they are not able to track it. So, management decide to allocate $25000 for upgrading their current tracking & product coding All the operation cost associated for manufacturing are given below. Now company is confused either to use RFID or Barcode or combined. They have options to use Active or passive RFID & barcodes. M1 M2 Labor Utilities Mat. Cost$3.00/box Company can produce 600 box in 10 days &each box contains 10 drugs & management want each drug to be monitored & they only want one reader for whole process $2000/month $2000/month $2.00/box - 2000/month Type Active RFID Barcode Passive RFID Cost/each 1.5 0.5 Reader $2500 $3000 $2500 Using the cost analysis identify which method they should implement. Consider the analysis for a 30 day time period.
Answer #1

Drug produced by the company in 10 days = 600*10*10 =60000

So, Drug produced by the company in 30 days = 60000*3 = 180000

As we know that total cost = Fixed cost + Variable Cost

We will first analyse all the options to arrive our conclusion

Option 1: 1.5/unit + $2500

TC =1.5*180000 + $2500 =$272500

Option 2: =0.5/unit + $3000

TC =0.5*180000 + $3000 = $ 93000

Option 3: =1/unit + $ 2500

TC =1* 180000 + $ 2500 = $182500

So option 2 is giving lesser TC thatn option 1 aand option 3. so, company can opt option 2.